Which of the following statements about gross profit is correct?

Multiple Choice

  • Gross profit = Net sales − Cost of goods sold.

  • Gross profit is recorded by a credit to the Gross Profit account.

  • A company sells $10,000 of goods. If the gross profit percentage is 32%, net income would be $3,200.

  • If net sales are $100 and cost of goods sold is $50 then the gross profit percentage is 100%.
Answer #1

Answer: Gross Profit = Net Sales -Cost of Goods Sold

Option 1 is Correct: Gross Profit is the Difference Between the Net sales Revenue and the Cost of Goods sold.

IF the Gross Profit is 32% so Gross Margin will be(10,000*32%)= 3,200 Not the net profit

Gross Margin = 50/100 = 50%not the 100%